# Hartman Effect from a Geometrodynamic Extension of Bohmian Mechanics

### v2

@@ 7 Concluding remarks
In this article, a solution to the quantum tunneling problem was constructed based on a geometrodynamic approach to Bohmian mechanics [37], which explains the Hartman effect [1] as a consequence of the distortion of spacetime generated by the quantum potential inside the barrier. To achieve this, the mathematical expressions for key quantities in the Bohmian framework, such as the quantum potential, momentum, and particle position, were determined from this solution. Subsequently, the mathematical expression for tunneling time in terms of the barrier width and the velocity of the ‘warp’ bubble was determined from the position expression.
From this time expression, it becomes evident that, as a result of the freedom in choosing the velocity $v_{s}$ caused by the spacetime’s own structure, hyperluminal particle tunneling is possible. Furthermore, it follows from this fact that by choosing a more general expression for $v_{s}$, the case in which there is saturation in tunneling times, as reported by Hartman [1], is not excluded either. However, the solution constructed in this article still reconciles the apparent contradiction between quantum mechanics and general relativity by considering a spacetime that allows for superluminal travel without violating the foundations of relativistic theory [38, 48]. It maintains the validity of explaining the Hartman effect as a consequence of the spacetime distortion generated by the quantum potential.
Finally, it is important to highlight that the solution constructed in this work, while acceptable as a possible explanation of the Hartman effect, also comes with the violation of energy conditions (when creating the ‘warp’ bubble inside the potential barrier). This would imply the need for exotic matter with unknown physical characteristics [38].
@@ Acknowledgments
This work was supported by the Coordination for the Improvement of Higher Education Personnel (CAPES) under Grant No. 88887.630121/2021-00, by the National Council for Scientific and Technological Development (CNPq) under Grants No. 309862/2021-3, No. 409673/2022-6, and No. 421792/2022-1, and by the National Institute for the Science and Technology of Quantum Information (INCT-IQ) under Grant No. 465469/2014-0.
Author contributions statement
S.L. and J.M. conceived the project. S.L. carried out the calculations under the supervision of J.M.. S.L. wrote the first version of the manuscript. All authors discussed and reviewed the manuscript.
8 Appendix: Solution of the differential equations (38), (56), (57), and (58)
In this appendix, the solution to the differential equations (38), (56), (57), and (58) is presented. This includes changes of variables, some algebraic procedures, and methods used for the resolution of each of these differential equations [49, 50, 51]. These expressions have a structure that seemingly leads to complicated differential equations. For this reason, and with the intention of facilitating the reading of this work, we will divide this appendix into three parts. The first part focuses on the solution of Eq. (38), the second part deals with the solution of Eqs. (56) and (57), while the third part concentrates on the solution of Eq. (58).
8.1 Part I
@@
where the constants $\alpha_{0}$, $\beta_{0}$, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\mu_{0}$, and $\mu_{1}$ were presented in Sec. 4 and are obtained by writing $[\alpha_{1}+\cos{(x)}]/[(\alpha_{1}+\cos{(x)})^{2}-\alpha_{0}^{2}\cos^{2}{(x)}]=\beta_{0}[z^{2}+\beta_{1}]/[z^{4}+2\beta_{3}z^{2}+\beta_{2}^{2}]$ from the Weierstrass substitution. Consequently, by integrating Eq. (76), we find
$\begin{split}S(z)=\frac{\alpha_{0}\beta_{0}}{2\sigma\sqrt{\beta_{3}^{2}-\beta_{2}^{2}}}\biggl{\{}\frac{\beta_{1}-\mu_{0}}{\sqrt{\mu_{0}}}\tan^{-1}{\left[\frac{z}{\sqrt{\mu_{0}}}\right]}-\frac{\beta_{1}-\mu_{1}}{\sqrt{\mu_{1}}}\tan^{-1}{\left[\frac{z}{\sqrt{\mu_{1}}}\right]}\biggr{\}},\end{split}$ (77)
@@ 8.2 Part II
$\begin{split}dU(z)=\frac{2\vartheta_{10}}{1-\vartheta_{12}}\biggl{\{}[1-\vartheta_{11}]\frac{dz}{1+z^{2}}+[\vartheta_{11}-\vartheta_{12}]\frac{dz}{\vartheta_{12}+z^{2}}\biggr{\}},\end{split}$ (82)
@@
Then, by integrating Eq. (82),
$\begin{split}U(\Xi_{I})=\frac{2\vartheta_{10}}{1-\vartheta_{12}}\biggl{\{}\frac{[1-\vartheta_{11}]}{2}\Xi_{I}+\frac{[\vartheta_{11}-\vartheta_{12}]}{\sqrt{\vartheta_{12}}}\tan^{-1}{\biggl{[}\frac{1}{\sqrt{\vartheta_{12}}}\tan{\biggl{(}\frac{\Xi_{I}}{2}\biggr{)}}\biggr{]}}\biggr{\}}+U_{0},\end{split}$ (83)
from which the general solution of the exact version of Eq. (56) is obtained by determining the value of $U_{0}(T)$. Therefore, by calculating $\partial_{T}U$ and equating it with the expression $\mu Q$, it follows that $U_{0}(T)=0$, as it is not possible to obtain a function that depends only on $T$ from the previous calculation. On the other hand, to determine the general solution from Eq. (57), it suffices to replace $\vartheta_{1}\rightarrow\vartheta_{5}$, $\vartheta_{2}\rightarrow\vartheta_{6}$, $\vartheta_{3}^{\prime}\rightarrow\vartheta_{7}$, $\vartheta_{4}\rightarrow\vartheta_{8}$, and $\Xi_{I}\rightarrow\Xi_{III}$ in each of the constants and in the angle of the general solution (83).
8.3 Part III
In a manner similar to the first part, equation (58) is reformulated as follows:
@@ 8.3 Part III
with $j=1,2,3,4,5,6$ and $l=1,2,3,4,5$. Then, following a procedure similar to that applied in Part I, we consider $2\sigma r_{s}=2\sigma(x-x_{s}(t))=X-T\equiv r$, allowing us to calculate the partial derivatives:
@@
clearly indicating that the differential equation (84) is not exact. Therefore, after calculating the integrating factor [49, 51]
@@
equation (84) transforms into an exact equation, where
@@
Nevertheless, finding the primitive of the above expression is not straightforward. However, if we recall that $r\equiv r(X,T)=2\sigma r_{s}(x,t)$ and also that $\sigma\ll 1$, we can calculate the second-order expansion [51] of the function $\mu(r)P(r)$ around the point $\mathcal{P}^{\prime}(0,0)$ given by
$-\frac{\iota_{j}^{\prime}\cosh^{j}{(r)}+1}{\iota_{j}^{\prime\prime}\cosh^{j}{(r)}-1}=\varsigma^{0}+\frac{\varsigma^{1}}{2}r^{2}.$ (91)
Here, the equality between the expansion and the function $\mu(r)P(r)$ is valid in this region $II$ within the context of small $\sigma$ parameters. Thus, by substituting Eq. (91) into Eq. (90) and integrating, we obtain
$\begin{split}U(r)=\{2\sigma\}^{-1}\{\varsigma^{0}r+\frac{\varsigma^{1}}{6}r^{3}\}+c(t)=\{2\sigma\}^{-1}\{\varsigma^{0}r+\frac{\varsigma^{1}}{6}r^{3}\}.\end{split}$ (92)
By calculating the derivative $\partial_{x}U$ and equating it with the expression $\mu(r)Q(r)$, we verify that Eq. (92) is necessary for $c(t)=0$, as happened in the case presented in Part $II$.

### v3

@@ 7 Concluding remarks
This work presents a geometrodynamic extension of Bohmian mechanics to explain quantum tunneling, where particles traverse barriers along geodesics in an Alcubierre-type spacetime. The derived model reveals a dual behavior in tunneling dynamics: for narrow barriers, the tunneling time depends on the barrier width, while for wide barriers, it saturates to a constant value—reproducing the Hartman effect. This saturation arises from a geometric self-regulation mechanism, where the quantum potential dynamically adjusts spacetime distortion to maintain a fixed tunneling time, consistent with relativistic causality despite effective superluminal propagation.
The framework bridges quantum mechanics and general relativity by attributing superluminal tunneling to local metric distortions rather than non-local effects or approximations. Crucially, it provides a new theoretical foundation to address longstanding paradoxes, particularly in light of recent experimental evidence, which challenges traditional Bohmian predictions (e.g., infinite dwell times in reflective barriers). By linking the Hartman effect to the dynamics of Alcubierre-like spacetime, our model resolves the saturation paradox without ad hoc assumptions, while preserving the causal structure of relativity.
Future directions for this research could explore experimental validation of the model’s predictions in systems where spacetime curvature and quantum tunneling naturally interplay—such as optomechanical cavities or graphene heterostructures—while also investigating connections to geometric phases like the Berry phase in topological materials, where curvature-induced effects may mimic the proposed geometrodynamic tunneling. Although the model relies on negative energy densities (analogous to the Casimir effect), recent advances in metamaterials and quantum field theory, including squeezed vacuum states and hyperbolic metamaterials, suggest promising avenues to engineer these conditions in laboratory settings.
This work not only advances the understanding of spacetime’s role in quantum phenomena but also underscores the need for beyond-standard interpretations to reconcile theory with emerging experimental frontiers. The geometrodynamic approach, though demanding in its mathematical and physical prerequisites, offers a viable pathway to unify quantum non-locality and relativistic constraints, paving the way for tests of quantum-gravity analogies in controlled settings.
@@ Acknowledgments
My dearest, first and foremost, I, SL, thank God the Father, infinite goodness, God the Son, redeemer of the world, and God the Holy Spirit, together with the Virgin Mary and Saint Joseph, for their protection, inspiration, and strength on this journey in pursuit of knowledge. I also thank my family, whose unwavering support shaped the person I am today.
This work was made possible through funding from the National Council for Scientific and Technological Development (CNPq), including the grant from the Institutional Training Program (PCI), Process No. 301066/2025-6, awarded by the Ministry of Science, Technology and Innovation (MCTI). Additional support came from the Coordination for the Improvement of Higher Education Personnel (CAPES) under Grant No. 88887.630121/2021-00, as well as CNPq Grants No. 309862/2021-3, No. 409673/2022-6, and No. 421792/2022-1. We also acknowledge the National Institute for the Science and Technology of Quantum Information (INCT-IQ) under Grant No. 465469/2014-0.
To all those who contributed directly or indirectly, we express our deepest gratitude.
Appendix A Solution of the differential equations (36), (55), (56), and (57)
In this appendix, the solution to the differential equations (36), (55), (56), and (57) is presented. This includes changes of variables, some algebraic procedures, and methods used for the resolution of each of these differential equations [53, 54, 55]. These expressions have a structure that seemingly leads to complicated differential equations. For this reason, and with the intention of facilitating the reading of this work, we divide this appendix into three parts. The first part focuses on the solution of Eq. (36), the second part deals with the solution of Eqs. (55) and (56), while the third part concentrates on the solution of Eq. (57).
A.1 Solution of Eq. (36)
@@ Appendix A Solution of the differential equations (36), (55), (56), and (57)
where the constants $\alpha_{0}$, $\beta_{0}$, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\mu_{0}$, and $\mu_{1}$ were presented in Sec. 4 and are obtained by writing $\alpha_{1}+\cos{(x)}]/[(\alpha_{1}+\cos{(x)})^{2}-\alpha_{0}^{2}\cos^{2}{(x)}]=\beta_{0}[z^{2}+\beta_{1}]/[z^{4}+2\beta_{3}z^{2}+\beta_{2}^{2}]$ from the Weierstrass substitution. Consequently, by integrating Eq. (86), we find
$S(z)$ $=$ $\frac{\alpha_{0}\beta_{0}}{2\sigma\sqrt{\beta_{3}^{2}-\beta_{2}^{2}}}\biggl\{\frac{\beta_{1}-\mu_{0}}{\sqrt{\mu_{0}}}\tan^{-1}{\left[\frac{z}{\sqrt{\mu_{0}}}\right]}$
- $\frac{\beta_{1}-\mu_{1}}{\sqrt{\mu_{1}}}\tan^{-1}{\left[\frac{z}{\sqrt{\mu_{1}}}\right]}\biggr\},$
@@
$dU(z)=\frac{2\vartheta_{10}}{1-\vartheta_{12}}\biggl\{[1-\vartheta_{11}]\frac{dz}{1+z^{2}}+[\vartheta_{11}-\vartheta_{12}]\frac{dz}{\vartheta_{12}+z^{2}}\biggr\},$ (92)
@@
Then, by integrating Eq. (92), we obtain
$U(\Xi_{I})$ $=$ $\frac{2\vartheta_{10}}{1-\vartheta_{12}}\biggl\{\frac{[1-\vartheta_{11}]}{2}\Xi_{I}+\frac{[\vartheta_{11}-\vartheta_{12}]}{\sqrt{\vartheta_{12}}}$
$\times$ $\tan^{-1}{\biggl[\frac{1}{\sqrt{\vartheta_{12}}}\tan{\biggl(\frac{\Xi_{I}}{2}\biggr)}\biggr]}\biggr\}+U_{0},$
from which the general solution of the exact version of Eq. (55) is obtained by determining the value of $U_{0}(T)$. Therefore, by calculating $\partial_{T}U$ and equating it with the expression $\mu Q$, it follows that $U_{0}(T)=0$, as it is not possible to obtain a function that depends only on $T$ from the previous calculation. On the other hand, to determine the general solution from Eq. (56), it suffices to replace $\vartheta_{1}\rightarrow\vartheta_{5}$, $\vartheta_{2}\rightarrow\vartheta_{6}$, $\vartheta_{3}^{\prime}\rightarrow\vartheta_{7}$, $\vartheta_{4}\rightarrow\vartheta_{8}$, and $\Xi_{I}\rightarrow\Xi_{III}$ in each of the constants and in the angle of the general solution (A.2).
A.3 solution of Eq. (57)
In a manner similar to that used in A.2, equation (57) is reformulated as follows:
@@
with $j=1,2,3,4,5,6$ and $l=1,2,3,4,5$. Then, following a procedure similar to that applied in A.2, we consider $2\sigma r_{s}=2\sigma(x-x_{s}(t))=X-T\equiv r$, allowing us to calculate the partial derivatives:
@@
clearly indicating that the differential equation (94) is not exact. Therefore, after calculating the integration factor [53, 55]
@@
Eq. (94) transforms into an exact equation, where
@@
The exact primitive of (100) is analytically challenging, but a systematic approximation can be developed for the small-parameter regime ($\sigma\ll 1$). Recalling that $r\equiv r(X,T)=2\sigma r_{s}(x,t)$ depends on both spatial and temporal coordinates, we perform a second-order bivariate Taylor expansion [55] of $\mu(r)P(r)=-[\iota_{j}^{\prime}\cosh^{j}{(r)}+1]/[\iota_{j}^{\prime\prime}\cosh^{j}{(r)}-1]$ around $\mathcal{P}^{\prime}(0,0)$:
$\mu(r)P(r)=\underbrace{\left.\mu P\right|_{(0,0)}}_{\varsigma^{0}}+\underbrace{\left.\frac{\partial(\mu P)}{\partial X}\right|_{(0,0)}\!\!\!X+\left.\frac{\partial(\mu P)}{\partial T}\right|_{(0,0)}\!\!\!T}_{\text{1st-order terms}}$
$+$ $\underbrace{\frac{1}{2}\left.\frac{\partial^{2}(\mu P)}{\partial X^{2}}\right|_{(0,0)}\!\!\!X^{2}+\left.\frac{\partial^{2}(\mu P)}{\partial X\partial T}\right|_{(0,0)}\!\!\!XT+\frac{1}{2}\left.\frac{\partial^{2}(\mu P)}{\partial T^{2}}\right|_{(0,0)}\!\!\!T^{2}}_{\text{2nd-order terms}}$
$+$ $\mathcal{O}(X^{3},X^{2}T,XT^{2},T^{3}),$
where:
  • 1.
    The zeroth-order term $\varsigma^{0}=\left.\mu P\right|_{(0,0)}$ captures the static background,
  • 2.
    The first-order derivatives vanish at $\mathcal{P}^{\prime}(0,0)$,
  • 3.
    The second-order coefficients govern the main nonlinear correction.
Higher-order terms ($\mathcal{O}(X^{3},X^{2}T,XT^{2},T^{3})$) are negligible under the low perturbation condition $\sigma\ll 1$, reducing the expansion to $\mu(r)P(r)\approx\varsigma^{0}+\varsigma^{1}r^{2}/2$. Substituting this into (100) and integrating yields to:
$U(r)$ $=$ $\{2\sigma\}^{-1}\{\varsigma^{0}r+\frac{\varsigma^{1}}{6}r^{3}\}+c(t)$
$=$ $\{2\sigma\}^{-1}\{\varsigma^{0}r+\frac{\varsigma^{1}}{6}r^{3}\}.$
where the integration constant $c(t)=0$ is required for consistency with $\partial_{x}U=\mu(r)Q(r)$, as in the analogous procedures shown in A.2.